OpenStudy (anonymous):

if two dice are thrown, what is the probability that the sum of the two will be 7?

OpenStudy (anonymous):

there is nothing that can go into 7 but itself

OpenStudy (anonymous):

so would it be 1/1?

OpenStudy (anonymous):

There are 36 different possibilities for how two dice can come up. How many of them are 7?

OpenStudy (anonymous):

But you can add more than one value to get 7. If you are throwing a six-sided die, then you have the possibility of getting 1, 2, 3, 4, 5, or 6 on either die.

OpenStudy (anonymous):

So you must consider the possibilities of getting a 1 and a 6. 1 + 6 = 7, right? So that could be a possibility. But you can also think of 3+4, 5+2, etc. And then you put those possibilities over the total possibilities.

OpenStudy (anonymous):

There are 36 different possible outcomes when rolling two dice. If you want the sum to be 7, then you must roll one of the following: (1,6), (2,5),(3,4), (4,3), (5,2), or (6,1) .6 possiblities So probability is 6/36 = 1/6

OpenStudy (anonymous):

hmmm...thanks

OpenStudy (anonymous):

You're welcome.